# Federated Learning: A new frontier in the exploration of multi-institutional medical imaging data

### v1

@@ Abstract
Artificial intelligence has transformed the perspective of medical imaging, leading to a genuine technological revolution in modern computer-assisted healthcare systems. However, ubiquitously featured deep learning (DL) systems require access to a considerable amount of data, facilitating proper knowledge extraction and generalization. Admission to such extensive resources may be hindered due to the time and effort required to convey ethical agreements, set up and carry the acquisition procedures through, and manage the datasets adequately with a particular emphasis on proper anonymization. One of the pivotal challenges in the DL field is data integration from various sources acquired using different hardware vendors, diverse acquisition protocols, experimental setups, and even inter-operator variabilities. In this paper, we review the federated learning (FL) concept that fosters the integration of large-scale heterogeneous datasets from multiple institutions in training DL models. In contrast to a centralized approach, the decentralized FL procedure promotes training DL models while preserving data privacy at each institution involved. We formulate the FL principle and comprehensively review general and dedicated medical imaging aggregation and learning algorithms, enabling the generation of a globally generalized model. We meticulously go through the challenges in constructing FL-based systems, such as data heterogeneity across the institutions, resilience to potential attacks on data privacy, and the variability in computational and communication resources among the entangled sites that might induce efficiency issues of the entire system. Finally, we explore the up-to-date open frameworks for rapid FL-based algorithm prototyping and shed light on future directions in this intensively growing field.
@@
federated learning, deep learning, heterogeneous data, multi-institutions, aggregation methods, malicious clients, privacy protection.
journal: arXiv.org
@@ 1 Introduction
Medical imaging (MI) is an intrinsic part of modern healthcare systems and plays an indispensable role in clinical and medical applications nowadays. The versatility of medical image analysis systems and their advancement over the last decades have enabled to inspect, measure and visualize the structure and function of different organs, allowing for early detection, precise diagnosis and prognosis, and meticulous therapy planning [Duncan and Ayache, 2000, Bankman, 2008, Frangi et al., 2023]. From initial screening to treatment selection and further follow-up, MI is critical in the healthcare chain. A variety of MI techniques and their derivatives founded on different phenomena, such as X-ray imaging, computed tomography (CT), magnetic resonance imaging (MRI) and positron emission tomography (PET), provide comprehensive and complementary information on macro- and microstructure [Lerch et al., 2017, Alexander et al., 2019, Novikov et al., 2019], as well as functional [Frangi et al., 2001, Deffieux et al., 2021], physiological [Cohen et al., 2017] and molecular characteristics [Gambhir, 2002], resulting in a better understanding of natural processes that occur in the body, including the maturation and aging leading to normative trajectories [Cox et al., 2016, Bethlehem et al., 2022], detection and progression of miscellaneous diseases [Basu et al., 2009, Mott et al., 2012, Chen, 2019, Planche et al., 2022], exploration the similarities between the species [de Schotten et al., 2019, Pagani et al., 2023] or discovery the genetic relationships with organ structure and/or conditions [Grasby et al., 2020, Stauffer et al., 2023].
@@
Over the last decade, we have been observing an authentic technological shift in medical image analysis and image-guided systems towards machine learning (ML) and predominantly deep learning (DL) techniques, leading to faster and more accurate data reconstruction, processing and inference [Shen et al., 2017, Litjens et al., 2017, Barragán-Montero et al., 2021]. DL-based procedures have already demonstrated extraordinary results in a broad spectrum of imaging protocols and tasks [Sahiner et al., 2019, Wang et al., 2020a, Ozturk et al., 2020, Wen et al., 2020, Arabi et al., 2021, Isensee et al., 2021, Sermesant et al., 2021], and are considered as the future of robotic-assisted surgeries [Li et al., 2022a, Varghese et al., 2024]. In order to improve the patient’s comfort and amenities as well as support clinicians, it is vital to explore the possibilities for optimizing MI phases with DL, including acquisition protocols, data reconstruction and processing procedures, and automation in making decisions that can significantly assist clinicians in everyday practice [Barragán-Montero et al., 2021]. As the DL system can integrate the data and improve inference by learning from large-scale observations, training and validating such models requires an extensive amount of diverse and labeled datasets to achieve high accuracy of the models [Fan et al., 2016, Bakas et al., 2018, Simpson et al., 2019, Littlejohns et al., 2020]. These datasets can be collected from multifarious sources, including diverse machine vendors, acquisition protocols, experimental setups, and the preprocessing pipelines employed. Gathering such broad-sided comprehensive datasets is challenging, mainly due to sensitive medical information that requires proper protection [Nelson, 2015, Rieke et al., 2020, Olatunji et al., 2024]. Although meticulous agreements are critical to ensure secure data sharing and adequate use by the community, the datasets must be compiled suitably, including anonymization, which goes beyond merely removing patient-identifying information and may include some interference with the raw data, such as face masking [Nelson, 2015, Olatunji et al., 2024, Willemink et al., 2020]. Therefore, it is unlikely that large and diverse datasets will be available shortly free of charge due to the considerable effort and time required for their collection, labeling and management [Rieke et al., 2020].
@@
Federated learning (FL) is a concept that addresses the need to integrate multi-institutional heterogeneous data in training DL models. FL is a decentralized approach, which means that DL models are trained and validated locally without sharing private data between the involved sites [McMahan et al., 2017, Rieke et al., 2020, Sheller et al., 2020, Li et al., 2020a]. In contrast to a traditional centralized procedure that operates on the data collected from different institutions on a single centralized server, in the FL-based approach, only the characteristics of locally trained models are exported from local collaborators to generate a global model (see Fig. 1). Developing an FL-based architecture requires carefully selected aggregation and learning strategies for a particular imaging modality and/or DL problem. The aggregation algorithm combines the parameters from locally trained models to yield a single global model [Qi et al., 2024a]. The quality of the new-created model vastly depends on the aggregation strategy chosen [Reddi et al., 2021, Li et al., 2020a, Wicaksana et al., 2023, Hosseini et al., 2023, Kalapaaking et al., 2023, Wu et al., 2023b].
@@
However, implementing the FL-based architecture demands further technical advancements to ensure optimal accuracy and versatility without compromising data security. In recent years, there has been an immense interest growth in federated techniques among the MI community in federated techniques (see Fig. 2), leading to the development of several unique procedures intended to address specific MI limitations and constraints, such as data heterogeneity [Li et al., 2020b, Feng et al., 2023, Zhao and Huang, 2023] or label deficiency [Wu et al., 2021, Peng et al., 2023].
@@ 1.1 Related works
Several review articles covering the FL concept in the context of MI have been published so far. Some of them demonstrate the application perspective in a targeted scope, including clinical oncology and cancer research [Chowdhury et al., 2021, Ankolekar et al., 2024, Bechar et al., 2025], radiomics [Guzzo et al., 2023], ocular imaging [Nguyen et al., 2022d], or radiology in general [Rehman et al., 2023]. In other reviews, the FL has been exposed as a part of the study with particular emphasis on a specific problem considered in MI such as brain tumor segmentation [Ahamed et al., 2023], diagnosis of common retinal diseases [Lim et al., 2024], COVID-19 detection [Mondal et al., 2023] or skin cancer screening [Wei et al., 2024]. MI has also been explored as one of the elements of a broader discussion covering the FL-based systems from the Smart Healthcare perspective, such as the Internet of Medical Things [Nguyen et al., 2022b, Chaddad et al., 2023, Shiranthika et al., 2023]. Alternatively, these reviews address specific FL-related challenges, e.g., the security and privacy preservation of medical data [Kaissis et al., 2020, Aouedi et al., 2022, Koutsoubis et al., 2024], uncertainty estimation [Koutsoubis et al., 2024] or lack of large sets of labeled data [Ng et al., 2021]. Other reports present the computational platforms and tools allowing for collaborative data analysis or rapid prototyping of the FL-based systems [Rootes-Murdy et al., 2022, Riviera et al., 2023].
Figure 2: The number of screened research papers on applications of federated learning in medical imaging per year. The results were taken from the MEDLINE and Google Scholar databases and followed the procedure presented in section 2.1. Data is up to date as of October 2024.
Only a few reports summarize the applications and challenges of FL in MI area without presenting strict mathematical formulations of aggregation and learning methods [Darzidehkalani et al., 2022, da Silva et al., 2023, Sandhu et al., 2023, Nazir and Kaleem, 2023, Guan et al., 2024]. Instead, these reviews present a general summary of FL methods without categorizing them by their dedicated problem-solving capabilities or concentrate solely on the applications of standard FL-based aggregation and learning algorithms to various imaging modalities, body organs, or DL tasks.
Search query MEDLINE Google Scholar Total w/o
duplicates
“federated learning” and “medical imaging” 68 99 147
“federated learning” and “neuroimaging” 25 98 113
“federated learning” and “MRI” 40 100 125
“federated learning” and “magnetic resonance imaging” 41 100 127
“federated learning” and “CT” 49 100 138
“federated learning” and “computed tomography” 26 100 119
“federated learning” and “tomography” 36 100 127
“federated learning” and “PET” 10 98 101
“federated learning” and “positron emission tomography” 9 99 102
“federated learning” and “x-ray” 41 99 125
“federated learning” and (“histopathology” or “histopathological”) 15 99 101
“federated learning” and “retinal” 12 100 105
“federated learning” and “OCT” 99 100 193
“federated learning” and “optical coherence tomography” 4 99 99
“federated learning” and “ultrasound” 11 100 105
“federated learning” and “mammography” 3 99 99
“federated learning” and “microscopic” 0 99 99
Table 1: The list of queries used in MEDLINE and Google Scholar engines to retrieve the preliminary list of research articles related to federated learning in medical imaging. The results from Google Scholar were limited to the first 100 entries from each query.
@@ 1.2 Contributions
The contributions of this review paper can be summarized as follows:
@@
    it provides a comprehensive overview of the theoretical aspects of FL, including recent advances in architectures, aggregation and learning approaches, and fundamental techniques used in privacy protection for local data in MI,
@@
    it gives an up-to-date overview of FL applications in the field of MI, with a particular emphasis on different imaging modalities and tasks,
@@
    it discusses open issues in FL and recent practices used to address them,
@@
    it provides an overview of open-source tools and frameworks created to facilitate the rapid prototyping of FL-based algorithms,
@@
    finally, it defines still unsolved challenges in FL and possible directions for further development.
@@ 1.3 Paper organization
The paper has been organized as follows. In the next section, we define our methodology for collecting the literature used in the study and illustrate the search results from existing databases, focusing on miscellaneous imaging modalities and DL applications. Section 3 covers the fundamental concepts of FL, starting with its definition and then smoothly introducing the basic optimization and aggregation techniques, the learning techniques, and the most common data privacy protection methods. Then, section 4 delves deeper into FL, presenting various topics and problems that must be adequately clarified before developing a practical architecture. In sections 5 and 6, we present an in-depth analysis of aggregation and learning methods used in MI. Section 7 reviews actively developed open-source FL frameworks for rapid prototyping and deployment. Finally, section 8 discusses the challenges of FL and proposes strategies to address them, guiding future developments in the field.
Figure 4: A. Distribution of imaging modalities and B. deep learning tasks in the examined federated learning research papers ($n=129$).
@@ 2.1 Search strategy
The initial search returned 2125 papers, of which 1086 were screened after removing duplicates. Among these, 479 articles did not meet the inclusion criteria based on their titles and abstracts. Two papers were not written in English, and the remaining articles were unrelated to FL applications in MI (447) or were Ph.D./Master theses (30). In the next round of screening, we excluded another 454 articles. These papers were either short conference abstracts, unavailable, or did not present advances in aggregation or learning techniques. A total of 153 articles were considered for this study, of which 24 were review papers included in Section 1.1. The remaining records are the research papers discussed in the main body of this article. Fig. 3 illustrates the flow chart of the screening procedure used in our examination.

### v2

@@ Abstract
Artificial intelligence has transformed the perspective of medical imaging, leading to a genuine technological revolution in modern computer-assisted healthcare systems. However, ubiquitously featured deep learning (DL) systems require access to a considerable amount of data, facilitating proper knowledge extraction and generalization. Access to such extensive resources may be hindered due to the time and effort required to convey ethical agreements, set up and carry the acquisition procedures through, and manage the datasets adequately with a particular emphasis on proper anonymization. One of the pivotal challenges in the DL field is data integration from various sources acquired using different hardware vendors, diverse acquisition protocols, experimental setups, and even inter-operator variabilities. In this paper, we review the federated learning (FL) concept that fosters the integration of large-scale heterogeneous datasets from multiple institutions in training DL models. In contrast to a centralized approach, the decentralized FL procedure promotes training DL models while preserving data privacy at each institution involved. We formulate the FL principle and comprehensively review general and specialized medical imaging aggregation and learning algorithms, enabling the generation of a globally generalized model. We meticulously go through the challenges in constructing FL-based systems, such as data and model heterogeneities across the institutions, resilience to potential attacks on data privacy, and the variability in computational and communication resources among the entangled sites that might induce efficiency issues of the entire system. Finally, we explore the up-to-date open frameworks for rapid FL-based algorithm prototyping, comprehensively present real-world implementations of FL systems and shed light on future directions in this intensively growing field.
@@
federated learning, deep learning, heterogeneous data, model heterogeneity, multi-institutions, aggregation methods, malicious clients, privacy protection.
journal: https://arxiv.org/
@@ 1 Introduction
Medical imaging (MI) is an intrinsic part of modern healthcare systems and plays an indispensable role in clinical and medical applications nowadays. The versatility of medical image analysis systems and their advancement over the last decades have enabled to inspect, measure and visualize the structure and function of different organs, allowing for early detection, precise diagnosis and prognosis, and meticulous therapy planning [Duncan and Ayache, 2000, Bankman, 2008, Frangi et al., 2023]. A variety of MI techniques and their derivatives founded on different phenomena, such as X-ray imaging, computed tomography (CT), magnetic resonance imaging (MRI) and positron emission tomography (PET), provide comprehensive and complementary information on macro- and microstructure [Lerch et al., 2017, Alexander et al., 2019, Novikov et al., 2019], as well as functional [Frangi et al., 2001, Deffieux et al., 2021], physiological [Cohen et al., 2017] and molecular characteristics [Gambhir, 2002], resulting in a better understanding of natural processes that occur in the body [Cox et al., 2016, Chen, 2019, Bethlehem et al., 2022].
@@
Over the last decade, we have been observing an authentic technological shift in medical image analysis and image-guided systems towards machine learning (ML) and predominantly deep learning (DL) techniques, leading to faster and more accurate data reconstruction, processing and inference [Shen et al., 2017, Litjens et al., 2017, Barragán-Montero et al., 2021]. DL-based procedures have already demonstrated extraordinary results in a broad spectrum of imaging protocols and tasks [Sahiner et al., 2019, Wang et al., 2020a, Ozturk et al., 2020, Wen et al., 2020, Arabi et al., 2021, Isensee et al., 2021, Sermesant et al., 2021], and are considered as the future of robotic-assisted surgeries [Li et al., 2022a, Varghese et al., 2024]. As the DL system can integrate data and improve inference by learning from large-scale observations, training and validating such models requires an extensive amount of diverse and labeled datasets [Fan et al., 2016, Bakas et al., 2018, Simpson et al., 2019, Littlejohns et al., 2020]. These datasets can be collected from multifarious sources, including diverse machine vendors, acquisition protocols, experimental setups, and the preprocessing pipelines employed. Gathering such broad-sided comprehensive datasets is challenging, mainly due to sensitive medical information that requires proper protection [Nelson, 2015, Rieke et al., 2020, Olatunji et al., 2024]. Although meticulous agreements are critical to ensure secure data sharing and adequate use by the community, the datasets must be compiled suitably, including anonymization, which goes beyond merely removing patient-identifying information and may include some interference with the raw data, such as face masking [Nelson, 2015, Olatunji et al., 2024, Willemink et al., 2020]. Therefore, it is unlikely that large and diverse datasets will be available shortly free of charge due to the considerable effort and time required for their collection, labeling and management [Rieke et al., 2020].
@@
Federated learning (FL) is a concept that addresses the need to integrate multi-institutional heterogeneous data in training DL models. FL is a decentralized approach, which means that DL models are trained and validated locally without sharing private data between the involved sites [McMahan et al., 2017, Rieke et al., 2020, Sheller et al., 2020, Li et al., 2020a]. In contrast to a traditional centralized procedure that operates on the data collected from different institutions on a single server, in the FL-based approach, only the characteristics of locally trained models are exported from local collaborators to generate a global model (see Fig. 1). Developing an FL-based architecture requires carefully selected aggregation and learning strategies for a particular imaging modality and/or DL problem. The aggregation algorithm combines the parameters from locally trained models to yield a single global model [Qi et al., 2024a]. The quality of the new-created model vastly depends on the aggregation strategy chosen [Reddi et al., 2021, Li et al., 2020a, Wicaksana et al., 2023, Hosseini et al., 2023, Kalapaaking et al., 2023, Wu et al., 2024].
@@
However, implementing the FL-based architecture demands further technical advancements to ensure optimal accuracy and versatility without compromising data security. In recent years, there has been immense growth in interest in federated techniques among the MI community (see Fig. 2), leading to the development of several unique procedures intended to address specific MI limitations and constraints, such as data heterogeneity [Li et al., 2020b, Feng et al., 2023, Zhao and Huang, 2023], model heterogeneity [Weng et al., 2025, Nezhad et al., 2025] or label deficiency [Wu et al., 2021, Peng et al., 2023].
@@ 1.1 Related works
Several review articles covering the FL concept in the context of MI have been published so far, as summarized in Table 1. Some of them demonstrate the application perspective in a targeted scope, including clinical oncology and cancer research [Ankolekar et al., 2025, Bechar et al., 2025], radiomics [Guzzo et al., 2023], ocular imaging [Nguyen et al., 2022c, Lim et al., 2024], or radiology in general [Rehman et al., 2023]. In other reviews, the FL has been exposed as a part of the study with particular emphasis on a specific problem considered in MI such as brain tumor segmentation [Ahamed et al., 2023], diagnosis of common retinal diseases [Lim et al., 2024], or COVID-19 detection [Mondal et al., 2023]. Alternatively, these reviews address specific FL-related challenges, e.g., the security and privacy preservation of medical data [Kaissis et al., 2020, Koutsoubis et al., 2024], uncertainty estimation [Koutsoubis et al., 2024] or lack of large sets of labeled data [Ng et al., 2021]. Other reports present the computational platforms and tools allowing for collaborative data analysis or rapid prototyping of the FL-based systems [Rootes-Murdy et al., 2022, Riviera et al., 2023].
Figure 2: The number of screened research articles on applications of federated learning in medical imaging per year. The results were obtained from the MEDLINE and Google Scholar databases and followed the procedure presented in section 2.1. Data is up to date as of November 2025.
Only a few reports summarize the applications and challenges of FL in MI area without presenting strict mathematical formulations of aggregation and learning methods [da Silva et al., 2023, Sandhu et al., 2023, Nazir and Kaleem, 2023, Guan et al., 2024, Bechar et al., 2025, Raza et al., 2025]. Instead, these reviews present a general summary of FL methods without categorizing them by their dedicated problem-solving capabilities or concentrate solely on the applications of standard FL-based aggregation and learning algorithms to various imaging modalities, body organs, or DL tasks.
@@ 1.2 Contributions
Our contributions can be summarized as follows:
@@
    we provide a comprehensive overview of the theoretical aspects of FL, including recent advances in architectures, aggregation and learning approaches, and fundamental techniques used in privacy protection for local data in MI,
@@
    we give an up-to-date overview of FL applications in the field of MI, with a particular emphasis on different imaging modalities and tasks,
@@
    we analyze distributions of aggregation strategies and learning methods used, categorized by imaging modality and deep learning tasks,
@@
    we discuss open issues in FL and recent practices used to address them, particularly in the context of real-world applications,
@@
    provide an issue–method–effect analysis that illustrates how different aggregation strategies and learning methods address key challenges in federated learning,
  6. 6.
    we provide an overview of open-source tools and frameworks created to facilitate the rapid prototyping of FL-based algorithms,
  7. 7.
    finally, we define still unsolved challenges in FL and possible directions for further development.
@@ 1.3 Paper organization
In the next section, we define our methodology for collecting the literature used in the study and illustrate the search results from existing databases, focusing on miscellaneous imaging modalities and DL applications. Section 3 covers the fundamental concepts of FL, starting with its definition and then smoothly introducing the basic optimization and aggregation techniques, the learning techniques, and the most common data privacy protection methods. Then, section 4 delves deeper into FL, presenting various topics and problems that must be adequately clarified before developing a practical architecture. In sections 5 and 6, we present an in-depth analysis of aggregation and learning methods used in MI. Section 7 reviews actively developed open-source FL frameworks, while section 8 presents examples of deploying federated learning in clinical settings. Finally, section 9 discusses the challenges of FL and proposes strategies to address them, guiding future developments in the field.
@@ 2.1 Search strategy
The initial search returned 2765 papers, of which 1901 were screened after removing duplicates. Among these, 887 articles did not meet the inclusion criteria based on their titles and abstracts. Seven papers were not written in English, and the remaining articles were unrelated to FL applications in MI (839) or were Ph.D./fig/Master theses (41). In the next round of screening, we excluded another 840 articles. These papers were either short conference abstracts, unavailable, did not present advances in aggregation or learning techniques, or reviews that were not focused on MI or FL. A total of 174 articles were considered for this study, of which 20 were review papers included in Section 1.1. The remaining records are the research papers discussed in the main body of this article. Fig. 3 illustrates the flow chart of the screening procedure used in our examination.
Reference MI modality Mathematical Dedicated to Technical FL challenges FL challenges FL frameworks Real-world
coverage principles MI algorithms presentation presentation solutions presentation examples
Kaissis et al. [2020] Multiple Narrow
Ng et al. [2021] MRI Descriptive Comprehensive
Darzidehkalani et al. [2022a] Multiple
Nguyen et al. [2022c] OCT,
retinal images
Descriptive Comprehensive
Rootes-Murdy et al. [2022] MRI
Ahamed et al. [2023] MRI
Guzzo et al. [2023] Multiple Descriptive Comprehensive
Mondal et al. [2023] Multiple Descriptive Comprehensive
Nazir and Kaleem [2023] Multiple Descriptive Comprehensive
Rehman et al. [2023] Multiple Detailed Comprehensive
Riviera et al. [2023] Comprehensive
Sandhu et al. [2023] Multiple Descriptive Comprehensive
da Silva et al. [2023] Multiple Descriptive Comprehensive
Sohan and Basalamah [2023] Multiple Comprehensive
Guan et al. [2024] Multiple Detailed Comprehensive
Koutsoubis et al. [2024] Multiple Detailed Narrow Mentioned
Lim et al. [2024] OCT,
retinal images
Comprehensive
Ankolekar et al. [2025] Multiple Descriptive Comprehensive
Bechar et al. [2025] Multiple Descriptive Comprehensive
Raza et al. [2025] Multiple Descriptive Comprehensive
This review Multiple Detailed Comprehensive
Table 1: Summary of review articles on the application of federated learning in medical imaging. Columns explanations: (2) MI modality coverage – number of medical imaging modalities discussed (multiple – more than two); (3) Mathematical principles – does the article coverage mathematical and algorithmic foundations of federated learning?; (4) Dedicated to MI algorithms – does the article focus on federated learning methods tailored to medical imaging?; (5) Technical presentation – level of technical presentation (descriptive – high-level discussion without mathematical detail, detailed – covers mathematical formulations of the methods); (6) FL challenges presentation – scope and depth of discussion of federated learning challenges such as data heterogeneity or malicious clients (narrow – superficial discussion on challenges only, comprehensive – coverage of multiple challenges in detail); (7) FL challenges solutions – does the article present discussion on the methods addressing specific federated learning challenges?; (8) FL frameworks presentation – does the article present the frameworks supporting the implementation of federated learning?; (9) Real-world examples – does the article discuss real-world deployment of federated learning in clinics? (mentioned – examples are only mentioned without details).
Search query MEDLINE Google Scholar Total w/o
duplicates
“federated learning” and “medical imaging” 153 100 226
“federated learning” and “neuroimaging” 38 100 126
“federated learning” and “MRI” 84 98 165
“federated learning” and “magnetic resonance imaging” 88 95 165
“federated learning” and “CT” 79 100 165
“federated learning” and “computed tomography” 46 100 135
“federated learning” and “tomography” 66 100 151
“federated learning” and “PET” 19 100 116
“federated learning” and “positron emission tomography” 12 100 108
“federated learning” and “x-ray” 61 100 144
“federated learning” and (“histopathology” or “histopathological”) 22 100 113
“federated learning” and “retinal” 25 100 115
“federated learning” and “OCT” 235 100 235
“federated learning” and “optical coherence tomography” 13 100 106
“federated learning” and “ultrasound” 23 100 122
“federated learning” and “mammography” 6 100 102
“federated learning” and “microscopic” 2 100 100
Table 2: The list of queries used in MEDLINE and Google Scholar engines to retrieve the preliminary list of research articles related to federated learning in medical imaging. The results from Google Scholar were limited to the first 100 entries from each query. Data is up to date as of November 2025.
Figure 4: A. Distribution of imaging modalities and B. deep learning tasks in the examined federated learning research papers in this review ($n=154$).
Figure 5: A,B Distribution of aggregation strategies and C,D learning methods, categorized by imaging modality (left diagrams) and deep learning task (right diagrams) across $n=154$ FL-based research articles considered in this study.